# Autoregressive Action Sequence Learning for Robotic Manipulation

### v4

@@ Abstract
Designing a universal policy architecture that performs well across diverse robots and task configurations remains a key challenge. In this work, we address this by representing robot actions as sequential data and generating actions through autoregressive sequence modeling. Existing autoregressive architectures generate end-effector waypoints sequentially as word tokens in language modeling, which are limited to low-frequency control tasks. Unlike language, robot actions are heterogeneous and often include high-frequency continuous values—such as joint positions, 2D pixel coordinates, and end-effector poses—which are not easily suited for language-based modeling. Based on this insight, we extend causal transformers’ single-token prediction to support predicting a variable number of tokens in a single step through our Chunking Causal Transformer (CCT). This enhancement enables robust performance across diverse tasks of various control frequencies, greater efficiency by having fewer autoregression steps, and lead to a hybrid action sequence design by mixing different types of actions and using a different chunk size for each action type. Based on CCT, we propose the Autoregressive Policy (ARP) architecture, which solves manipulation tasks by generating hybrid action sequences. We evaluate ARP across diverse robotic manipulation environments, including Push-T, ALOHA, and RLBench, and show that ARP, as a universal architecture, matches or outperforms the environment-specific state-of-the-art in all tested benchmarks, while being more efficient in computation and parameter sizes. Videos of our real robot demonstrations, all source code and the pretrained models of ARP can be found at http://github.com/mlzxy/arp.
I Introduction
@@
Autoregressive models are the basis of recent breakthroughs in natural language processing [1]. These models predict the next token in a sequence based on the previous tokens. Autoregressive models are typically implemented as causal transformers, where each token attends only to preceding ones, and they are trained with the single objective of maximizing the conditional likelihood of each token. Despite their simplicity, autoregressive models such as GPTs [2] are shown to demonstrate a reasoning ability that can capture causal dependencies [3]. In this work, we present a new universal autoregressive architecture that can be used for various robot manipulation tasks in diverse environments.
@@
Decision Transformer (DT) and Trajectory Transformer (TT) are two pioneering approaches that use autoregressive models to solve control tasks [4, 5]. These methods learn to generate trajectories as $(R_{1},s_{1},a_{1},\dots,R_{T},s_{T},a_{T})$, where $R_{t},s_{t},a_{t}$ respectively denote the reward-to-go [6], the state, and the action at time-step $t$. However, these methods are primarily applied to tasks with fully observed, low-dimensional states—which is rarely the case in robotics applications. Recent work focuses on applying autoregression only on action sequences, such as Gato [7], VIMA [8], ManipLLM [9]. Despite their impressive results, these methods remain limited to low-frequency control tasks, as they represent robot actions with key end-effector waypoints and generate one action at a time, similar to word generation in language modeling.
@@
However, unlike language, robot actions are heterogeneous and include continuous values—such as joint positions, 2D pixel coordinates, and effector poses. Additionally, in high-frequency control tasks, continuous actions are expected to maintain temporal smoothness—a requirement absent in language modeling. To adapt autoregressive models for robotic tasks, we propose the Chunking Causal Transformer (CCT), an improved version of the causal transformer used in standard autoregressive models. CCT introduces an important modification: it predicts the future tokens (a chunk of actions) from empty tokens rather than from the original sequence, as illustrated in Figure 6. In doing so, CCT extends the next-single-token prediction of causal transformer to chunking autoregression—next-multi-token prediction in a single step. Despite its simplicity, chunking autoregression offers three key advantages:
@@
    Chunking autoregression increases efficiency by reducing the number of inference passes that are required.
@@
While action chunking has already been introduced in the Action Chunking Transformer (ACT) by [10], ACT is a one-step prediction model with a fixed chunk size. Instead, our approach supports variable chunk sizes for generating hybrid action sequences. Figure 10 shows that our method outperforms ACT by a significant margin in all environments. Figure 9 shows that our chunking autoregression is the key factor behind this strong performance. We illustrate the essential difference between action chunking, standard autoregression, and our chunking autoregression in Figure 1.
@@
Figure 1: Existing Works versus Our Autoregressive Policy. Action chunking models (left) predict all action tokens in a single step [10, 11, 12]. Standard autoregression models (middle) generate one action token in each step, which is inefficient and unsuitable for high-frequency control tasks [4, 5, 7, 8, 9]. Our proposed chunking autoregression (right) generates a chunk of variable number of action tokens per step, offering greater efficiency, strong performance across diverse tasks, and flexibility in designing hybrid action sequences. We compare the performance of these three action prediction strategies in Figure 9. Note all models use Model Predictive Control to predict $L$ actions, execute them, update the observation, and then predict actions again. Autoregressive generation is performed without executing actions or changing the current observation.
@@
To summarize, our contributions are threefold. (1) We propose the Chunking Causal Transformer (CCT), which extends the single-token prediction of causal transformer to multi-token prediction, and therefore enables chunking autoregression. We also design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing, as shown in Figure 6. (2) Based on our CCT, we present the Auto-regressive Policy (ARP), a model that learns to generate heterogeneous action sequences autoregressively for solving robotic manipulation tasks. The ARP architecture is summarized in Figure 4. (3) We evaluate the same ARP architecture across Push-T [13], ALOHA [10], and RLBench [14], three environments with diverse manipulation tasks and control modalities, as outlined in Figure 2. Our study shows that ARP matches or outperforms all environment-specific SoTAs, while being more efficient computationally and using smaller parameter sizes, as in Figure 8. In addition, we evaluate ARP with a real robot on a challenging, contact-rich nut-tightening task, as in Figure 12.
@@
Figure 2: Overview of the simulation environments. We evaluate our method on Push-T, ALOHA, and RLBench, three task suites with significantly different properties and requirements. Push-T [13] requires many steps to complete (long horizon) and where same sub-goals can be reached in various ways (multi-modality). ALOHA [10] has a high-dimensional action space (14 joints of two robot arms), a high control frequency (50Hz), and a short time limit (8 secs). RLBench [14] has only the gripper pose as action but contains 18 different language-conditioned tasks.
II Related Work
@@
Learning robotic manipulation from demonstrations. Imitation learning enables robots to learn to perform tasks demonstrated by experts [15]. Recently, various methods have been developed for manipulation learning with different task constraints and control modalities. Notably, [13] proposed the Diffusion Policy (DP) method for solving the Push-T task. [10] proposed the Action Chunking Transformer (ACT) for bi-manual manipulation tasks in the ALOHA environment. [16] proposed RVT-2 for language-conditioned tasks in the RLBench environment [14]. We outline these environments and the corresponding state-of-the-art (SoTA) solutions in Figure 2 and Figure A1, respectively. In contrast, our proposed autoregressive policy is a universal architecture that outperforms each environment-specific SoTA on Push-T, ALOHA, and RLBench.
@@
Autoregressive models for control tasks. Besides the pioneering Decision Transformer and Trajectory Transformer, recent works such as VIMA [8], Gato [7], GR1 [12], OpenVLA [11] and ManipLLM [9] have looked into designing autoregressive models for robotic tasks. Despite the impressive results, these approaches are limited to low-frequency control tasks that rely on end-effector waypoints [11]. GR1 and OpenVLA only use autoregression in their LLM backbones and do not apply autoregression to solve control tasks. Most of the existing works require fine-tuning a large language model (LLM) such as LLaMA [9] to include target end-effector poses within text-based responses or predict poses from LLM’s hidden features. The reliance on resource-intensive LLMs leads to large computational overhead, even for tasks that could be addressed with lightweight models. Without these constraints, our autoregressive policy outperforms SoTAs in multiple environments while being more efficient in computation and parameter sizes.
Figure 3: Learned Action Sequences. In Push-T, our model predicts a sequence of high-level waypoints, followed by a sequence of low-level positions that connect the waypoints together and form the pushing trajectory, analogous to hierarchical planning [17]. In ALOHA, we predict the joint values and then end-effector waypoints conditioned on the joint values, a process akin to forward kinematics [18]. We bypass the waypoint generation during inference. In RLBench, we predict the target end-effector’s position first, then gripper rotation and state in that position. For our real robot experiment, we define a set of primitive actions, as detailed in section IV-C. We predict the action type and then continuous values of that action.
@@
Hierarchical policies. Planning actions on multiple levels of abstraction is an important ability [19]. Existing methods generally separate the designs of low-level and high-level policies, and uses different modules for the different levels of abstraction [19, 20]. This complicated procedure prohibits a wider application. In contrast, our autoregressive policy is able to capture the underlying causal dependencies in robotic tasks by predicting a sequence of actions of different levels of abstraction by using a single model, as shown by our action sequence designs for diverse environments in Figure 3.
III Method
In this section, we present the Auto-regressive Policy (ARP), built upon our Chunking Causal Transformer (CCT), which generates robot action sequences autoregressively. We summarize the architecture in Figure 4.
@@ III Method
Figure 4: Autoregressive Policy Architecture. A sequence of past actions and a chunk of empty tokens are concatenated and projected into embeddings. Empty tokens correspond to future actions, which are unknown and need to be predicted. These embeddings are fed into our Chunking Causal Transformer (CCT) along with the vision features of the current observation. CCT alternates between self-attention within the sequence embeddings and cross-attention with the vision features. Self-attention is causal for the input actions and bidirectional among the empty tokens. Distributions of future actions are decoded from the updated embeddings of the empty tokens.
@@
Embedding and decoding heterogeneous actions. Language models map each word to a continuous vector called word embedding. The word embeddings of the input sentences are fed into a causal transformer. The distribution of the next word is decoded from the output embedding of the last word with a linear layer. Figure 5 and 7 illustrate our embedding and decoding methods for robot actions. Discrete actions are embedded by a table lookup on a weight matrix and decoded into a categorical distribution with a linear layer, similar to words in language modeling. Continuous actions are embedded with a linear layer and decoded into the parameters of a Gaussian mixture distribution with another linear layer. Actions that are defined as pixel coordinates are embedded by retrieving the point-wise features at the coordinates on a visual feature map. The output spatial distribution is obtained by multiplying the output embedding with the visual feature map, and converting the result into a 2-d heatmap with the up-sampling operator from RAFT [21].
Figure 5: Embeddings for Discrete, Continuous, and Pixel-coordinate Actions. Discrete actions are embedded by a simple table lookup on a weight matrix. Continuous actions are embedded with a linear layer. Pixel-coordinate actions are embedded by retrieving the point-wise features at the coordinates on the visual feature maps.
Figure 6: Training Chunking Causal Transformer (CCT) with Teacher-forcing. Causal transformers are trained efficiently with only a single forward pass for all tokens in a given sequence. However, suppose $a_{2},a_{3}$ and $a_{4}$ are in separate chunks, the CCT forward passes of predicting $a_{2},a_{3}$ and $a_{4}$ cannot be merged directly. Naively running separate passes significantly increases computation costs, as in Figure A4. With the proposed attention interleaving, we can update all the empty tokens in a single pass, regardless of the number of tokens or variable chunk size configurations. The key idea is to update empty tokens and action tokens separately and reuse the causally attended action tokens to update empty tokens. A step-by-step example is provided in Figure A7 and Video/attention-interleaving-tour.mp4 in the supplementary.
@@
Chunking causal transformer. Figure A3 illustrates the essential difference between a causal transformer and our CCT. A causal transformer modifies the token embedding with causal attention so that the last token becomes the next token. Our CCT modifies the token embedding with causal attention for the action tokens $a_{i}$ and bidirectional attention for the empty tokens $e_{i}$ (future actions). The empty tokens become the next tokens. This allows the prediction of a chunk of variable number of next tokens at once in a single forward pass by adding empty tokens, which enables action chunking during autoregressive generation. We study the impacts of our chunking autoregression in detail in Section IV. In ARP, CCT alternates between self-attention within the input embeddings and cross-attention with vision features, as in Figure 4. We extract vision features from a standard backbone identical to the ones used in SoTA methods, as detailed in section IV.
@@
Train-time attention interleaving. During training, a causal transformer is taught to predict each token in a given sequence by consuming all previous ground-truth tokens as input. This training strategy is named teacher-forcing [22]. As shown in Figure 6, only a single forward pass is required for training samples such as $a_{1},a_{2},a_{3}\rightarrow a_{4}$ (predict $a_{4}$ from $a_{1},a_{2},a_{3}$), $a_{1},a_{2}\rightarrow a_{3}$, and $a_{1}\rightarrow a_{2}$. Causal transformers are therefore efficiently trained with teacher-forcing. We follow this teacher-forcing strategy. However, training CCT yields separate forward passes per chunk. For example, the prediction of $a_{4}$ depends on $a_{2},a_{3}$, as in $a_{1},a_{2},a_{3},e_{4}\rightarrow a_{4}$, but $a_{2},a_{3}$ need to be replaced with $e_{2},e_{3}$ to predict them from $a_{1}$, as in $a_{1},e_{2},e_{3}\rightarrow a_{2},a_{3}$. This prohibits the use of a single forward pass for both $a_{1},a_{2},a_{3},e_{4}\rightarrow a_{4}$ and $a_{1},e_{2},e_{3}\rightarrow a_{2},a_{3}$. Note $a_{i}$ denotes the $i$-th action and $e_{i}$ denotes the empty token for $i$-th action. This issue increases the training cost and complicates the training procedure.
@@
To resolve this, we have all action tokens and their corresponding empty tokens as model input, as in ➊ ➋ of Figure A7. We then compute bidirectional attention within empty tokens and causal attention from empty tokens to action tokens using two masked attentions, as in ➌ ➍. Next, we compute causal attention within action tokens, as in ➎. This leverages the fact that action tokens are independent of future empty tokens. As a result, the updated action tokens are computed once and reused in the next layer, as in ➏. This enables a single forward pass of all tokens in three attention operations, regardless of the number of tokens or variable chunk configurations. We name this procedure attention interleaving. Figure A4 demonstrates the reduced MACs of training with attention interleaving. We implement attention interleaving as an internal acceleration mechanism of the transformer layer, which is transparent to other network modules. Note that attention interleaving is only used during training and incurs no additional inference cost.
Figure 7: Decoders for Discrete, Continuous, and Pixel-coordinate Actions. For discrete actions, we decode the action embeddings into a categorical distribution with a linear layer followed by a softmax operation. For continuous actions, we decode the embeddings into the parameters of a Gaussian mixture distribution with a linear layer. For the pixel-coordinate actions, we multiply the embedding with a visual feature map or a weight tensor, and convert the result into a 2-d heatmap.
Environment Method Success Rate #MACs #Params
Push-T Diffusion Policy 78.8 6.8G 25.5M
ARP (Ours) 87.1 3.7G 23.5M
ALOHA ACT 20.8 80.8 17.8G 50.9M
ARP (Ours) 24.8 94 17.8G 47.6M
RLBench RVT-2 81.4 57.1G 72.1M
ARP (Ours) 81.6 56.2G 71.9M
84.9 57.4G 73.8M
Figure 8: Comparing our Autoregressive Policy to the SoTA of each environment. Our autoregressive policy (ARP) outperforms environment-specific SoTA and is more efficient in MACs (number of multiply-accumulates) and parameter sizes. We report results of the transformer version of the diffusion policy because of its overall better performance. The RVT-2 [16] results are reported from the original paper. All MACs and parameter sizes are measured with THOP [23]. We list the success rates of both the insertion (left) and cube transfer (right) tasks in Gym-ALOHA [24]. Per task success rates of RLBench are summarized in Figure A5.
@@
Inference. During the test rollouts, we extract vision tokens from the current observation and provide them as input to ARP, which then generates actions autoregressively by sampling from the decoded action distribution and appending the selected actions to the existing action sequence. This process of generating and executing actions is repeated until episode termination (success, failure, or reaching the time limit). Actions are generated according to the sequence formats shown in Figure 3. We manually set the chunk size for each type of action, and total sequence length for each task. We provide more implementation details and hyper-parameter values in Section IV and Appendix A-B.
IV Experiments

### v5

@@ Abstract
Designing a universal policy architecture that performs well across diverse robots and task configurations remains a key challenge. In this work, we address this by representing robot actions as sequential data and generating actions through autoregressive sequence modeling. Existing autoregressive architectures generate end-effector waypoints sequentially as word tokens in language modeling, which are limited to low-frequency control tasks. Unlike language, robot actions are heterogeneous and often include continuous values—such as joint positions, 2D pixel coordinates, and end-effector poses—which are not easily suited for language-based modeling. Based on this insight, we introduce a straightforward enhancement: we extend causal transformers’ single-token prediction to support predicting a variable number of tokens in a single step through our Chunking Causal Transformer (CCT). This enhancement enables robust performance across diverse tasks of various control frequencies, greater efficiency by having fewer autoregression steps, and lead to a hybrid action sequence design by mixing different types of actions and using a different chunk size for each action type. Based on CCT, we propose the Autoregressive Policy (ARP) architecture, which solves manipulation tasks by generating hybrid action sequences. We evaluate ARP across diverse robotic manipulation environments, including Push-T, ALOHA, and RLBench, and show that ARP, as a universal architecture, outperforms the environment-specific state-of-the-art in all tested benchmarks, while being more efficient in computation and parameter sizes. Videos of our real robot demonstrations, all source code and the pretrained models of ARP can be found at http://github.com/mlzxy/arp.
1 Introduction
@@
Autoregressive models are the basis of recent breakthroughs in natural language processing (min2023recent). These models predict the next token in a sequence based on the previous tokens. Autoregressive models are implemented with causal transformers, where each token attends only to preceding ones, and they are trained with the single objective of maximizing the conditional likelihood of each token. Despite their simplicity, autoregressive models such as GPTs (mann2020gpt3) are shown to demonstrate a reasoning ability that can capture causal dependencies (prystawski2024stepbystep). In this work, we explore the design of a universal autoregressive architecture that can be used for various robot manipulation tasks in diverse environments.
@@
Decision Transformer (DT) and Trajectory Transformer (TT) are the pioneering approaches that use autoregressive models to solve control tasks (chen2021decisionT; janner2021trajT). These methods learn to generate trajectories as $(R_{1},s_{1},a_{1},\dots,R_{T},s_{T},a_{T})$, where $R_{t},s_{t},a_{t}$ respectively denote the reward-to-go (tamar2016rewardtogo), the state, and the action at time-step $t$. However, these methods are primarily applied to tasks with fully observed, low-dimensional states—a condition rarely met in robotics applications. Recent work focuses on applying autoregression only on action sequences, such as Gato (reed2022gato), VIMA (jiang2022vima), ManipLLM (li2024manipllm). Despite the impressive results, these methods remain limited to low-frequency control tasks, as they represent robot actions with key end-effector waypoints and generate one action at a time, similar to word generation in language modeling.
@@
However, unlike language, robot actions are heterogeneous and include continuous values—such as joint positions, 2D pixel coordinates, and effector poses. Additionally, in high-frequency control tasks, continuous actions are expected to maintain temporal smoothness—a requirement absent in language modeling. To adapt autoregressive models for robotic tasks, we propose the Chunking Causal Transformer (CCT), an improved version of the causal transformer used in standard autoregressive models. CCT introduces a simple modification: it predicts the future tokens, or a chunk of actions, from empty tokens rather than from the original sequence, as illustrated in Figure 5. In doing so, CCT extends the next-single-token prediction of causal transformer to chunking autoregression—next-multi-token prediction in a single step. Despite the simplicity of our enhancement, chunking autoregression offers three key advantages:
@@
    Chunking autoregression increases efficiency by reducing the inference passes required.
@@
While action chunking has been previously introduced in the Action Chunking Transformer (ACT) by zhao2023aloha, ACT is a one-step prediction model with a fixed chunk size. Instead, our approach supports variable chunk sizes for more powerful autoregressive models in generating hybrid action sequences. Table 2 shows that our method outperforms ACT by a significant margin in all environments. Table 1 shows that our chunking autoregression is the key factor behind the strong performance of the proposed model. We illustrate the essential difference between action chunking, standard autoregression, and our chunking autoregression in Figure 1.
@@
Figure 1: Existing Works versus Our Autoregressive Policy. Action chunking models (left) predict all action tokens in a single step (zhao2023aloha; kim2024openvla; wu2023gr1). Standard autoregression models (middle) generate one action token in each step, which is inefficient and unsuitable for high-frequency control tasks chen2021decisionT; janner2021trajT; reed2022gato; jiang2022vima; li2024manipllm. Our proposed chunking autoregression (right) generates a chunk of variable number of action tokens per step, offering greater efficiency, strong performance across diverse tasks, and flexibility in designing hybrid action sequences. We compare the performance of these three action prediction strategies in Table 1. Note all models use Model Predictive Control to predict $L$ actions, execute them, update the observation, and then predict actions again. Autoregressive generation is performed without executing actions or changing the current observation.
@@
To summarize, our contributions are threefold. (1) We propose the Chunking Causal Transformer (CCT), which extends the single-token prediction of causal transformer to multi-token prediction, and therefore enables chunking autoregression. We also design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing, as shown in Figure 5. (2) Based on our CCT, we present the Auto-regressive Policy (ARP), a model that learns to generate heterogeneous action sequences autoregressively for solving robotic manipulation tasks. The ARP architecture is summarized in Figure 3. (3) We evaluate the same ARP architecture across Push-T (chi2023diffp), ALOHA (zhao2023aloha), and RLBench (james2020rlbench), three environments with diverse manipulation tasks and control modalities, as outlined in Figure 2. Our study shows that ARP outperforms all environment-specific SoTAs, while being more efficient computationally and using smaller parameter sizes, as in Figure 6. In addition, we evaluate ARP with a real robot on a challenging, contact-rich nut-tightening task, as shown in Figure 9.
@@
Figure 2: Overview of the simulation environments. We evaluate our technique on Push-T, ALOHA, and RLBench, three task suites with significantly different properties and requirements. Push-T (chi2023diffp) is a task that requires many steps to complete (long horizon) and where the same sub-goals can be reached in various ways (multi-modality). ALOHA (zhao2023aloha) has a high-dimensional action space (14 joints of two robot arms), a high control frequency (50Hz), and a short time limit (8 seconds). RLBench (james2020rlbench) has only the gripper pose as an action but contains 18 different language-conditioned tasks.
2 Related Work
@@
Learning robotic manipulation from demonstrations. Imitation learning enables robots to learn to perform tasks demonstrated by experts (zare2024ilsurvey). Recently, various methods have been developed for manipulation learning with different task constraints and control modalities. Notably, chi2023diffp proposed the Diffusion Policy (DP) method for solving the Push-T task. zhao2023aloha proposed the Action Chunking Transformer (ACT) for bi-manual manipulation tasks in the ALOHA environment. goyal2024rvt2 proposed RVT-2 for language-conditioned tasks in the RLBench environment (james2020rlbench). We outline these environments and the corresponding state-of-the-art (SoTA) solutions in Figure 2 and Figure 7, respectively. In contrast, our proposed autoregressive policy is a universal architecture that outperforms each environment-specific SoTA on Push-T, ALOHA, and RLBench.
@@
Autoregressive models for control tasks. Besides the pioneering Decision Transformer and Trajectory Transformer, recent works such as VIMA (jiang2022vima), Gato (reed2022gato), GR1 (wu2023gr1), OpenVLA (kim2024openvla) and ManipLLM (li2024manipllm) have looked into designing autoregressive models for robotic tasks. Despite the impressive results, these approaches are limited to low-frequency control tasks that rely on end-effector waypoints (kim2024openvla). GR1 and OpenVLA only use autoregression in their LLM backbones and do not apply autoregression to solve control tasks. Most of the existing works require fine-tuning a large language model (LLM) such as LLaMA (li2024manipllm) to include target end-effector poses within text-based responses or predict poses from LLM’s hidden features. The reliance on resource-intensive LLMs leads to large computational overhead, even for tasks that could be addressed with lightweight models. Without these constraints, our autoregressive policy outperforms SoTAs in multiple environments while being more efficient in computation and parameter sizes.
@@
Hierarchical policies. Planning actions on multiple levels of abstraction is an important ability (pateria2021hiera-survey). Existing methods generally separate the designs of low-level and high-level policies, and uses different modules for the different levels of abstraction (le2018hiera-imi-rl; pateria2021hiera-survey; belkhale2023hydra; belkhale2024rth; chen2024hiera-diffu). This complicated procedure prohibits a wider application. In contrast, our autoregressive policy is able to capture the underlying causal dependencies in robotic tasks by predicting a sequence of actions of different levels of abstraction by using a single model, as shown by our action sequence designs for diverse environments in Figure 4.
@@
Figure 3: Autoregressive Policy Architecture. A sequence of past actions and a chunk of empty tokens are concatenated and projected into embeddings. The empty tokens correspond to future actions, which are unknown and need to be predicted. These embeddings are fed into our Chunking Causal Transformer (CCT) along with the vision features of the current observation. CCT alternates between self-attention within the input embeddings and cross-attention with the vision features. Self-attention is causal for the input actions and bidirectional among the empty tokens. Distributions of future actions are decoded from the updated embeddings of the empty tokens.
3 Method
In this section, we present the Auto-regressive Policy (ARP), which predicts actions using the Chunking Causal Transformer (CCT). The architecture is summarized in Figure 3.
@@ 3 Method
Embedding and decoding heterogeneous actions. Language models map each word to a continuous vector called word embedding. The word embeddings of the input sentences are fed into a causal transformer. The distribution of the next word is decoded from the output embedding of the last word with a linear layer. Figure A2 and A3 illustrate our embedding and decoding methods for robot actions. Discrete actions are embedded by a table lookup on a weight matrix and decoded into a categorical distribution with a linear layer, similar to words in language modeling. Continuous actions are embedded with a linear layer and decoded into the parameters of a Gaussian mixture distribution with another linear layer. Actions that are defined as pixel coordinates are embedded by retrieving the pointwise features at the coordinates on a visual feature map. The output spatial distribution is obtained by multiplying the output embedding with the visual feature map, and converting the result into a 2d heatmap with the upsampling operator from RAFT (teed2020raft).
Figure 4: Learned Action Sequences. In Push-T, our model predicts a sequence of high-level waypoints, followed by a sequence of low-level positions that connect the waypoints together and form in the pushing trajectory, analogous to hierarchical planning (hafner2022hiera-plan). In ALOHA, we predict the joint values and then the end-effector waypoints conditioned on the joint values, a process akin to forward kinematics (kucuk2006robot). We bypass the waypoint generation during inference. In RLBench, we predict the target end-effector’s position first, then the gripper rotation and state in that position. For our real robot experiment, we define a set of primitive actions, as detailed in section 4.3. We predict the action type and then the continuous values of that action.
@@
Chunking causal transformer. Figure A1 illustrates the essential difference between a causal transformer and our CCT. A causal transformer modifies the token embedding with causal attention so that the last token becomes the next token. Our CCT modifies the token embedding with causal attention for the action tokens $a_{i}$ (past actions) and bidirectional attention for the empty tokens $e_{i}$ (future actions). The empty tokens become the next tokens. This allows the prediction of a chunk of variable number of next tokens at once in a single forward pass by adding empty tokens, which enables action chunking during autoregressive generation. We study the impacts of our chunking autoregression in detail in Section 4. In ARP, CCT alternates between self-attention within the input embeddings and cross-attention with vision features, as in Figure 3. We extract vision features from a standard backbone identical to the ones used in SoTA methods, as detailed in section 4.
@@
Train-time attention interleaving. During training, a causal transformer is taught to predict each token in a given sequence by consuming all previous ground-truth tokens as input. This training strategy is named teacher-forcing (teacher-forcing). As shown in Figure 5, only a single forward pass is required for training samples such as $a_{1},a_{2},a_{3}\rightarrow a_{4}$ (predict $a_{4}$ from $a_{1},a_{2},a_{3}$), $a_{1},a_{2}\rightarrow a_{3}$, and $a_{1}\rightarrow a_{2}$. Causal transformers are therefore efficiently trained with teacher-forcing. We follow this teacher-forcing strategy. However, training CCT yields separate forward passes per chunk. For example, the prediction of $a_{4}$ depends on $a_{2},a_{3}$, as in $a_{1},a_{2},a_{3},e_{4}\rightarrow a_{4}$, but $a_{2},a_{3}$ need to be replaced with $e_{2},e_{3}$ to predict them from $a_{1}$, as in $a_{1},e_{2},e_{3}\rightarrow a_{2},a_{3}$. This prohibits the use of a single forward pass for both $a_{1},a_{2},a_{3},e_{4}\rightarrow a_{4}$ and $a_{1},e_{2},e_{3}\rightarrow a_{2},a_{3}$. Note $a_{i}$ denotes the $i$-th action and $e_{i}$ denotes the empty token for $i$-th action. This issue increases the training cost and drastically complicates the training procedure.
@@
To resolve this, we first divide the attention into three categories: (1) causal attention within the action tokens, (2) bidirectional attention within the empty tokens, and (3) causal attention between the empty and the action tokens, as marked in Figure 5. Figure A4 shows that the causal attention within action tokens is computed repeatedly. We avoid this redundancy by precomputing the causal attentions for all action tokens and caching the results. For the empty tokens, we combine their inner bidirectional attention and the causal attention toward cached action tokens. This enables a single forward pass of all tokens in three attention operations, regardless of the number of tokens. We name this procedure attention interleaving. Figure A4 demonstrates the reduced MACs of training with attention interleaving. We implement attention interleaving as an internal acceleration mechanism of the transformer layer, which is transparent to other network modules. Note that attention interleaving is only used during training and incurs no additional inference cost.
@@
Inference. During the test rollouts, we extract vision tokens from the current observation and provide them as input to ARP, which then generates actions autoregressively by sampling from the decoded action distribution and appending the selected actions to the existing action sequence. This process of generating and executing actions is repeated until episode termination (success, failure, or reaching the time limit). Actions are generated according to the sequence formats shown in Figure 4. We manually set the chunk size for each type of action, and total sequence length for each task. We provide more implementation details and hyper-parameter values in Section 4 and Appendix A.2.
Figure 5: Training Chunking Causal Transformer (CCT) with Teacher-forcing. Causal transformers are trained efficiently with only a single forward pass for all tokens in a given sequence. However, suppose $a_{2},a_{3}$ and $a_{4}$ are in separate chunks, the CCT forward passes of predicting $a_{2},a_{3}$ and $a_{4}$ cannot be merged directly. Naively running separate passes significantly increases computation costs, as in Figure A4. With the proposed attention interleaving, we precompute and cache the causal attention results for all action tokens $a_{1},a_{2},a_{3}$. For the empty tokens $e_{2},e_{3},e_{4}$, we combine their bidirectional attention (lightgray) and the causal attention towards action tokens (green, blue). Since all the computations related to the action tokens are cached, we can update all the empty tokens in a single pass, regardless of the number of tokens. An example of attention interleaving is provided in the Video/attention-interleaving-tour.mp4 in the supplementary.
4 Experiments